OpenStudy (cucacula):

find dy/dx of the following functions : y=(3+x/3-x)^4 anyone?

OpenStudy (mertsj):

\[y=(\frac{3+x}{3-x})^4\]

OpenStudy (mertsj):

First use the power rule. Then use the quotient rule on the fractional part.

OpenStudy (cucacula):

need to find U and V first? @Mertsj

OpenStudy (e.mccormick):

Well, there would be a chain rule for the overall. The outer power is one part, and everything inside the paren is the other.

OpenStudy (cucacula):

so the U = 3 + x and V = 4 - x ? @e.mccormick

OpenStudy (e.mccormick):

Well, the part inside the paren can be done using the quotent rule, but the chain rule must be applied to the overall. They use u and v all over calc, so what is u and what is v depends on which rule you are talking about.

OpenStudy (anonymous):

\[y= (\frac{ 3 + x }{ 3 - x })^{4}\] can be written as \[y= \frac{ (3 + x)^{4} }{ (3 - x)^{4} }\]

OpenStudy (anonymous):

the quotient rule (differentiation) says that when you have \[y = \frac{ u }{ v }\] and you differentiate it, \[\frac{ dy }{ dx } = \frac{ (v \times \frac{ du }{ dx })-(u \times \frac{ dv }{ dx }) }{ v ^{2} }\]

OpenStudy (cucacula):

but need to find U and V again? @kausarsalley

OpenStudy (anonymous):